# Benchmarking Ultra-Low-Power $\mu$NPUs

### v1

@@ 7. ACKNOWLEDGMENTS
90 Model Stage MAX78k (C MAX78k (R) GAP8 NXP-MCXN947 HX-WE2 (S) HX-WE2 (P) STM32H7A3ZI ESP32 Time Power Time Power Time Power Time Power Time Power Time Power Time Power Time Power NAS Initialization 0.07±0.001 13.67±0.02 0.17±0.002 12.84±0.03 12.94±0.237 102.93±3.11 0.22±0.001 112.29±3.39 2.60±0.002 91.50±0.22 2.60±0.001 91.44±0.19 0.43±0.005 47.38±0.07 86.73±0.001 105.93±0.09 Memory I/O 21.24±0.004 25.05±0.05 35.53±0.003 22.66±0.03 2.66±0.007 38.58±0.95 0.05±0.001 117.26±0.84 0.12±0.001 97.55±1.25 0.12±0.001 97.01±0.71 Inference 4.63±0.002 80.41±0.08 4.64±0.002 79.03±0.05 77.40±0.470 117.22±1.47 27.27±0.002 118.03±0.65 8.99±0.022 112.35±0.76 8.32±0.008 114.62±0.76 550.01±0.012 54.91±0.09 468.38±0.001 151.99±0.31 Post-Processing 0.02±0.001 25.66±0.04 0.11±0.001 22.43±0.06 0.08±0.001 33.23±1.28 0.01±0.001 109.62±1.19 0.01±0.001 91.19±1.31 0.01±0.001 91.17±1.39 0.02±0.001 46.01±0.05 0.02±0.001 101.93±1.86 Idle - 13.21±0.02 - 10.87±0.01 - 33.67±0.35 - 105.71±0.14 - 89.09±1.30 - 89.09±1.30 - 38.13±0.05 - 77.73±0.07 ResNet Initialization 0.07±0.001 13.87±0.05 0.17±0.001 12.82±0.07 10.13±0.001 145.90±3.59 0.22±0.001 110.37±1.99 2.59±0.001 94.27±0.23 2.60±0.001 93.51±2.01 0.43±0.002 48.32±0.07 87.49±0.059 106.41±0.07 Memory I/O 26.53±0.002 25.14±0.05 44.87±0.003 22.64±0.04 2.47±0.110 38.39±0.31 0.05±0.001 115.95±0.99 0.12±0.001 99.15±0.45 0.14±0.001 97.74±1.37 Inference 2.89±0.001 47.12±0.09 2.96±0.001 46.86±0.07 38.45±0.081 113.47±0.40 16.07±0.002 118.57±0.52 8.30±0.001 110.13±0.79 8.54±0.015 109.71±2.35 282.00±0.010 54.98±0.08 318.15±0.001 144.91±0.27 Post-Processing 0.12±0.001 26.10±0.06 0.96±0.002 22.45±0.05 0.04±0.001 37.41±1.55 0.04±0.001 109.55±1.95 0.04±0.001 93.19±0.15 0.04±0.001 91.95±1.27 0.08±0.002 50.47±0.09 0.05±0.001 102.79±3.71 Idle - 13.21±0.02 - 10.87±0.01 - 33.67±0.35 - 105.71±0.14 - 89.09±1.30 - 89.09±1.30 - 38.13±0.05 - 77.73±0.07 SimpleNet Initialization 0.07±0.001 13.71±0.04 0.17±0.002 12.81±0.03 10.02±0.002 137.44±3.95 0.22±0.001 108.98±0.42 2.56±0.003 95.75±0.49 2.60±0.001 91.51±0.21 0.43±0.001 48.13±0.04 87.37±0.001 106.58±0.66 Memory I/O 12.14±0.003 24.46±0.07 20.22±0.003 22.27±0.04 2.60±0.076 38.51±0.69 0.05±0.001 117.54±1.69 0.12±0.001 91.11±1.77 0.12±0.001 96.97±0.69 Inference 2.68±0.002 50.89±0.04 2.67±0.003 49.86±0.05 38.23±0.117 122.12±3.70 13.52±0.001 118.40±0.81 8.10±0.001 107.15±0.69 7.13±0.001 106.84±0.77 276.03±0.011 55.19±0.06 283.21±0.002 139.96±1.84 Post-Processing 0.11±0.001 25.62±0.07 0.86±0.001 22.51±0.05 0.04±0.001 36.83±0.39 0.04±0.001 109.33±0.21 0.04±0.001 90.54±0.14 0.04±0.001 90.47±0.21 0.10±0.002 46.10±0.07 0.05±0.001 105.11±3.25 Idle - 13.21±0.02 - 10.87±0.01 - 33.67±0.35 - 105.71±0.14 - 89.09±1.30 - 89.09±1.30 - 38.13±0.05 - 77.73±0.07 AutoEncoder Initialization 0.07±0.001 13.75±0.06 0.17±0.001 12.86±0.05 8.17±0.076 89.22±1.28 0.28±0.001 109.14±0.20 2.60±0.002 90.79±0.70 2.60±0.001 96.04±1.01 0.12±0.002 48.17±0.06 19.76±0.002 103.51±0.59 Memory I/O 9.66±0.001 25.04±0.08 15.40±0.003 22.63±0.06 2.63±0.042 38.55±0.56 0.15±0.001 112.87±1.21 0.03±0.001 92.75±1.83 0.03±0.001 94.03±0.62 Inference 0.14±0.001 81.67±0.09 0.16±0.001 81.34±0.07 0.67±0.015 84.35±1.25 19.25±0.003 112.19±0.44 0.45±0.001 92.40±1.19 0.75±0.001 99.02±1.56 5.01±0.110 56.11±0.09 7.11±0.001 157.17±0.17 Post-Processing - - - - - - - - - - - - - - - - Idle - 13.21±0.02 - 10.87±0.01 - 33.67±0.35 - 105.71±0.14 - 89.09±1.30 - 89.09±1.30 - 38.13±0.05 - 77.73±0.07 YoloV1 Initialization 0.07±0.001 13.90±0.03 0.17±0.001 12.87±0.02 7.46±0.001 96.28±2.90 0.28±0.001 108.85±0.39 2.60±0.003 100.94±1.01 2.60±0.003 100.45±0.59 3.64±0.002 49.04±0.05 628.97±0.002 106.49±0.163 Memory I/O 8.84±0.003 41.82±0.04 11.99±0.002 35.55±0.04 2.83±0.037 38.10±0.11 0.51±0.001 110.51±1.61 1.11±0.001 100.31±3.09 1.11±0.001 101.95±2.90 Inference 2.61±0.002 21.47±0.04 2.62±0.001 21.13±0.05 36.11±0.047 113.67±2.39 19.25±0.001 113.31±1.66 8.48±0.001 112.63±1.30 7.30±0.001 111.69±0.97 336.00±0.013 54.99±0.08 536.22±0.001 129.74±0.107 Post-Processing 0.54±0.001 25.87±0.06 3.82±0.001 22.41±0.04 1.82±0.002 39.89±0.62 0.11±0.001 109.35±0.15 0.42±0.001 93.63±1.45 0.42±0.001 95.65±0.79 0.48±0.001 45.71±0.04 0.53±0.001 109.57±0.08 Idle - 13.21±0.02 - 10.87±0.01 - 33.67±0.35 - 105.71±0.14 - 89.09±1.30 - 89.09±1.30 - 38.13±0.05 - 77.74±0.07

### v2

@@ 7. ACKNOWLEDGMENTS
90 Model Stage MAX78k (C MAX78k (R) GAP8 NXP-MCXN947 HX-WE2 (S) HX-WE2 (P) STM32H7A3ZI ESP32 Time Power Time Power Time Power Time Power Time Power Time Power Time Power Time Power NAS Initialization 0.07±0.001 13.67±0.02 0.17±0.002 12.84±0.03 12.94±0.237 102.93±3.11 0.24±0.001 25.40±3.39 2.60±0.002 91.50±0.22 2.60±0.001 91.44±0.19 0.43±0.005 47.38±0.07 86.73±0.001 105.93±0.09 Memory I/O 21.24±0.004 25.05±0.05 35.53±0.003 22.66±0.03 2.66±0.007 38.58±0.95 0.83±0.001 24.52±0.84 0.12±0.001 97.55±1.25 0.12±0.001 97.01±0.71 Inference 4.63±0.002 80.41±0.08 4.64±0.002 79.03±0.05 77.40±0.470 117.22±1.47 24.86±0.002 36.38±0.65 8.99±0.022 112.35±0.76 8.32±0.008 114.62±0.76 550.01±0.012 54.91±0.09 468.38±0.001 151.99±0.31 Post-Processing 0.02±0.001 25.66±0.04 0.11±0.001 22.43±0.06 0.08±0.001 33.23±1.28 0.02±0.001 24.03±1.19 0.01±0.001 91.19±1.31 0.01±0.001 91.17±1.39 0.02±0.001 46.01±0.05 0.02±0.001 101.93±1.86 Idle - 13.21±0.02 - 10.87±0.01 - 33.67±0.35 - 23.53±0.14 - 89.09±1.30 - 89.09±1.30 - 38.13±0.05 - 77.73±0.07 ResNet Initialization 0.07±0.001 13.87±0.05 0.17±0.001 12.82±0.07 10.13±0.001 145.90±3.59 0.24±0.001 25.14±1.99 2.59±0.001 94.27±0.23 2.60±0.001 93.51±2.01 0.43±0.002 48.32±0.07 87.49±0.059 106.41±0.07 Memory I/O 26.53±0.002 25.14±0.05 44.87±0.003 22.64±0.04 2.47±0.110 38.39±0.31 0.91±0.001 24.89±0.99 0.12±0.001 99.15±0.45 0.14±0.001 97.74±1.37 Inference 2.89±0.001 47.12±0.09 2.96±0.001 46.86±0.07 38.45±0.081 113.47±0.40 12.97±0.002 36.69±0.52 8.30±0.001 110.13±0.79 8.54±0.015 109.71±2.35 282.00±0.010 54.98±0.08 318.15±0.001 144.91±0.27 Post-Processing 0.12±0.001 26.10±0.06 0.96±0.002 22.45±0.05 0.04±0.001 37.41±1.55 0.04±0.001 24.60±1.95 0.04±0.001 93.19±0.15 0.04±0.001 91.95±1.27 0.08±0.002 50.47±0.09 0.05±0.001 102.79±3.71 Idle - 13.21±0.02 - 10.87±0.01 - 33.67±0.35 - 23.62±0.14 - 89.09±1.30 - 89.09±1.30 - 38.13±0.05 - 77.73±0.07 SimpleNet Initialization 0.07±0.001 13.71±0.04 0.17±0.002 12.81±0.03 10.02±0.002 137.44±3.95 0.24±0.001 25.23±0.42 2.56±0.003 95.75±0.49 2.60±0.001 91.51±0.21 0.43±0.001 48.13±0.04 87.37±0.001 106.58±0.66 Memory I/O 12.14±0.003 24.46±0.07 20.22±0.003 22.27±0.04 2.60±0.076 38.51±0.69 0.91±0.001 24.84±1.69 0.12±0.001 91.11±1.77 0.12±0.001 96.97±0.69 Inference 2.68±0.002 50.89±0.04 2.67±0.003 49.86±0.05 38.23±0.117 122.12±3.70 12.49±0.001 35.61±0.81 8.10±0.001 107.15±0.69 7.13±0.001 106.84±0.77 276.03±0.011 55.19±0.06 283.21±0.002 139.96±1.84 Post-Processing 0.11±0.001 25.62±0.07 0.86±0.001 22.51±0.05 0.04±0.001 36.83±0.39 0.14±0.001 24.87±0.21 0.04±0.001 90.54±0.14 0.04±0.001 90.47±0.21 0.10±0.002 46.10±0.07 0.05±0.001 105.11±3.25 Idle - 13.21±0.02 - 10.87±0.01 - 33.67±0.35 - 23.73±0.14 - 89.09±1.30 - 89.09±1.30 - 38.13±0.05 - 77.73±0.07 Autoenc Initialization 0.07±0.001 13.75±0.06 0.17±0.001 12.86±0.05 8.17±0.076 89.22±1.28 0.23±0.001 24.95±0.20 2.60±0.002 90.79±0.70 2.60±0.001 96.04±1.01 0.12±0.002 48.17±0.06 19.76±0.002 103.51±0.59 Memory I/O 9.66±0.001 25.04±0.08 15.40±0.003 22.63±0.06 2.63±0.042 38.55±0.56 0.22±0.001 24.01±1.21 0.03±0.001 92.75±1.83 0.03±0.001 94.03±0.62 Inference 0.14±0.001 81.67±0.09 0.16±0.001 81.34±0.07 0.67±0.015 84.35±1.25 0.51±0.003 31.49±0.44 0.45±0.001 92.40±1.19 0.75±0.001 99.02±1.56 5.01±0.110 56.11±0.09 7.11±0.001 157.17±0.17 Post-Processing - - - - - - - - - - - - - - - - Idle - 13.21±0.02 - 10.87±0.01 - 33.67±0.35 - 23.53±0.14 - 89.09±1.30 - 89.09±1.30 - 38.13±0.05 - 77.73±0.07 YoloV1 Initialization 0.07±0.001 13.90±0.03 0.17±0.001 12.87±0.02 7.46±0.001 96.28±2.90 0.28±0.001 25.66±0.39 2.60±0.003 100.94±1.01 2.60±0.003 100.45±0.59 3.64±0.002 49.04±0.05 628.97±0.002 106.49±0.163 Memory I/O 8.84±0.003 41.82±0.04 11.99±0.002 35.55±0.04 2.83±0.037 38.10±0.11 0.77±0.001 25.41±1.61 1.11±0.001 100.31±3.09 1.11±0.001 101.95±2.90 Inference 2.61±0.002 21.47±0.04 2.62±0.001 21.13±0.05 36.11±0.047 113.67±2.39 17.18±0.001 28.43±1.66 8.48±0.001 112.63±1.30 7.30±0.001 111.69±0.97 336.00±0.013 54.99±0.08 536.22±0.001 129.74±0.107 Post-Processing 0.54±0.001 25.87±0.06 3.82±0.001 22.41±0.04 1.82±0.002 39.89±0.62 1.27±0.001 24.08±0.15 0.42±0.001 93.63±1.45 0.42±0.001 95.65±0.79 0.48±0.001 45.71±0.04 0.53±0.001 109.57±0.08 Idle - 13.21±0.02 - 10.87±0.01 - 33.67±0.35 - 22.91±0.14 - 89.09±1.30 - 89.09±1.30 - 38.13±0.05 - 77.74±0.07